please what is the dilution ratio of caterpillar force inceticide per 16 litre of water and at what area of land should solution covered
1
Expert's answer
2019-09-19T09:37:22-0400

For caterpillar force insecticide exists special dose (in liquid state): 0.35 ml per liter of water (300 liters for covering 1 ha). So, we can calculate the amount of insecticide for 16 liters of water: 16 * 0.35 = 5.6 liters.

Now let`s compute the dilution ratio for this substance: 5.6 / (5.6+16) = 25.9%

The solution should cover area (10000m2 * 16) / 300 = 533 m2